OpenStudy (rabekaromero):

Consider the following pair of equations: y = x + 4 y = −2x − 2 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

KjSaif (kjsaif):

@563blackghost

OpenStudy (joshmorris):

I think like this y= -2(x+4) -2

OpenStudy (rabekaromero):

Why would you put the first equation into the second one?

OpenStudy (joshmorris):

x+4 = –2x – 2 3x = -6 x = ?

OpenStudy (rabekaromero):

-2

OpenStudy (joshmorris):

Yes

OpenStudy (joshmorris):

x = -2 and y = 2 so you would write the solutions as you would an (x,y) point or (-2, 2)

OpenStudy (rabekaromero):

okay

OpenStudy (joshmorris):

Medal?

OpenStudy (joshmorris):

R u going to give me medal?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)